Accelerator Update for April 1, 2009 - April 3, 2009 |
Wednesday April 1 The day shift began with Operations monitoring the Tevatron (TeV) store 6950 (146.88E30) and the Recycler stash (186.9E10), with the Antiproton Source stacking (32.4E10), with the Linac off due to a quadrupole power supply problem, with the Booster, Main Injector (MI), and the MiniBooNE and NuMI experiments all waiting for Linac beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. The Linac resumed sending beam at 8:23 AM. At 12:21 PM, the MI Power Supply interlock Loop (MI-20U) tripped off. Operations contacted power supply technicians and the Operations MI Specialist. A power supply tech reported at 1:04 PM that the lambda power supply at MI-20 was back online. MI resumed sending beam. A LCW technician at CUB reported at 1:48 PM that he had put column one into internal circulation mode. At 3:06 PM, the Booster Debuncher Klystron PFN tripped off. It reset. Operations reported at 3:48 PM that the Recycler had stashed antiprotons six times during the day shift. Operations terminated store 6950 at 5 PM. Cryo system technicians began work on a TeV sector E4 wet engine at 5:12 PM. Operations prepared D0 and CDF for access at 5:20 PM. Cryo system techs reported at 7 PM that they had completed their wet engine work. Also, CDF came out of access. D0 came out of access at 7:13 PM. Operations put the TeV through a dry squeeze at 7:45 PM. Operations put the TeV through a wet squeeze at 8:39 PM. Operations put the TeV into shot setup at 9:02 PM. Operations established store 6953 at 10:22 PM with an initial luminosity of 322.92E30. Operations reported at 11:59 PM that the Recycler had stashed seven times during the evening shift. Thursday April 2 The midnight shift began with Operations monitoring store 6953 (224.15E30) and the stash (156.9E10), with Pbar stacking (21.2E10), and with NuMI and MiniBooNE taking beam. At 5:03 AM, a Booster vacuum pump (IP122) tripped off. It reset. Operations reported at 7:56 AM that the Recycler had stashed eight times during the midnight shift. At 8:08 AM, a Booster RF station (BRF13) tripped off. It reset. The Neutron Therapy Facility (NTF) reported at 8:30 AM that they would be conducting 2.5 hours of collimator transmission studies. Operations terminated store 6953 at 9:53 AM. Operations put the TeV into shot setup at 10:02 AM. Operations reported at 11:17 AM that LCW techs at CUB brought column one back online from its regeneration. Operations established store 6954 at 11:21 AM with an initial luminosity of 329.64E30. A MI Physicist began a study at 1:24 PM. Operations reported at 3:59 PM that the Recycler had stashed six times during the day shift. At 8:35 PM, LRF3 tripped off. It reset. At 10:34 PM, Operators noticed that a number of Booster correction magnets had their ramp tables zeroed. Further investigation uncovered a CAMAC crate with failed power supply. Booster resumed sending beam at 11:34 PM. Operators replaced the CAMAC crate. Operations reported at 11:57 PM that the Recycler had stashed eight times during the evening shift. Friday April 3 The midnight shift began with Operations monitoring store 6954 (93.0E30) and the stash (319.9E10), with Pbar stacking (26.7E10), and with NuMI and MiniBooNE taking beam. A MI Physicist reported at 12:32 AM that he had changed the horizontal collimation beam orbit. The collimators were using new position parameters, but the LVDT readbacks were not tracking as well as they should have. At 1:38 AM, LRF3 tripped off for the first of three times in three minutes. It reset. Operations terminated store 6954 at 2:49 AM. At 3:01 AM, KRF5 tripped off on a “coil comparator” indication. Operators contacted the Operations Proton Source Specialist and then consulted with the Linac on-call tech. Operators eventually replaced the coil power supply. Operations put the TeV into shot setup at 3:02 AM. Operations put Pbar into standby at 3:11 AM for the KRF5 work. Operators brought KRF5 back online at 4:20 AM. Operations established store 6955 at 5:45 AM with an initial luminosity of 307.3E30. NTF reported at 7:19 AM that they had begun setting up their beam line for a patient treatment later in the morning. Operations reported at 8:03 AM that the Recycler had stashed five times during the midnight shift. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 04/06/2009 email Fermilab |
| Security, Privacy, Legal | |